OpenStudy (anonymous):

An investor sold 100 shares of stock valued at $34.50 per share. The stock was purchased at $7.25 per share two years ago. Find the rate of continuously compounded interest that would be necessary in a banking account for the investor to make the same profit. I need to solve using the formula A=Pe^rt please help as this is extra credit and i really need it i have a bad grade i just need some guidance how to start this world problem

OpenStudy (anonymous):

A = 34.50*100; P = 7.25*100; t =2 A = Pe^rt or (A/P) = e^rt ln(A/P) = ln(e^rt) ln(A) - ln(P) = rt You should be able to solve for "r" from this equation

OpenStudy (anonymous):

Thankyou !

OpenStudy (anonymous):

you are welcome!

OpenStudy (anonymous):

If you have the answer, then you need to close the question.

OpenStudy (anonymous):

oh okay thanks will do that (: